Suppose that the price of a bottle of wine is €25 in Paris, and the wine is sold at $40 in New York. What will be the exchange rate if the law of one price holds?

please show steps
Answer #1

If the law of One price will be holding, It will mean that the price of one commodity will be similar in different nation and in such case the exchange rate would be as follows-

1 Euro= (40/25)= 1.60 Dollars

1 Dollar= (25/40)= .625 Euros.